A 8.40 kg rock whose density is 4900 kg/m3 is suspended by a string such that half of the rock's volume is under water. 1-What is the tension in the string?​
Answer #1

There are 3 forces acting

m*g acting downward

m*g = 8.40*9.8 = 82.32 N

volume of object = mass/density

              = 8.40/4900

              = 1.714*10^-3 m^3

volume insode water =( 1.714*10^-3)/2 = 8.57*10^-4 m^3

Buoyant force acting upward

FB = density of liquid * volume * g

= 1000 kg/m^3 * 8.57*10^-4 m^3 * 9.8 m/s^2

=8.4 N

Tension is acting upward

balance these 3 forces

T + FB = m*g

T + 8.4 = 82.32

T = 73.9 N

Answer: 73.9 N